OpenStudy (anonymous):

Change ((-3sqrt3)/2, (3/2)) to polar coordinates?

OpenStudy (anonymous):

R=x^2+y^2 Q=arc tan (y/x) (X,y)=(p,q)

OpenStudy (anonymous):

Sorry meant (r,q)

OpenStudy (anonymous):

ok so what do i plug in for all of that?

OpenStudy (anonymous):

Your values for x and y

OpenStudy (anonymous):

R^2= (-3sqrt3/2)^2+(3/2)^2= (-3*3/4)+(9/4)=-9/4+9/4=0 Arc tan (-3sqrt3/2/(3/2)=arc tan (-sqrt3)

OpenStudy (anonymous):

Q=2pi/3

OpenStudy (anonymous):

ok thanx so much!!!

OpenStudy (anonymous):

Answer (0,2pi/3)

OpenStudy (anonymous):

I was working it through and checked here to ask if it was right but you wrote the same thing i got!

OpenStudy (anonymous):

Awesome glad I could help